• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name J. Singh

    Log entry time 07:37:51 on July21,2004

    Entry number 129123

    keyword=TECHS: LHRS status

    LQ1:
    during the access, Peter and Ryan rest LQ1 by hand following S. Choi's instructions from 1.5 yrs ago. afterwards, i was able to successfully reset and power on the LQ1 power supply. i set the current to 20A and the power supply got to 20A. i then set the current to 1748.05787A and it slowly starting ramping. after a few min, the LHe level dipped again and LQ1 tripped off again. i called the guys downstairs to reset LQ1 again. at 0728, LQ1 is ramping up again, and the LHe looks more stable...we'll stay in controlled access until the LQ1 has safely ramped back up.

    LDipole:
    After waiting for about 1/2 hour after the LDipole crashed, i tried clearing the last remaining fault (M3/M4 Overtemp). it worked. i turned on the power supply and set the current to 20A. it worked. i changed the set momentum to the LHRS to 2.97100 and then quickly back to 2.97100. the left dipole is ramping up slowly to a set current of 1040.3353A. as of 0736, the left dipole is back up.


    FIGURE 1

    FIGURE 2